OpenStudy (anonymous):

Carla drew two triangles; triangle ABC and triangle PQR, on a grid. She planned to cut out the two triangles to make a flag. The vertices of triangle ABC are at A(1, 4), B(-2, -1), and C(3, -2). The vertices of triangle PQR are at P(-3, 0), Q(-2, -4), and R(2, 3). Using the coordinates of the vertices of each triangle explain whether the two triangles are congruent, similar or neither. Give me a sec and I will graph this

OpenStudy (anonymous):

There they are sketched out.

OpenStudy (anonymous):

They don't look congruent or similar.

OpenStudy (anonymous):

One is a right triangle and the other looks obtuse

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

None of the dimensions are evenly distributed between the 2 triangles either.

OpenStudy (apoorvk):

Great, okay! Now we don't go by looks in such cases, looks can be deceptive :P So what we need to do her is that, we find out all the sides of the 2 triangles, that is you will need to use the distance formula 6 times. And then, is all 3 pair of sides are equal, then they are congruent, if the ratios of side-lengths are constant and equal then they are similar triangles! And. If none - then none. -_-

OpenStudy (anonymous):

What is the distance formula?

OpenStudy (apoorvk):

Hmm For two points, say A(x\(_1\),y\(_1\)) and B(x\(_2\),y\(_2\)), the distance AB between them is: \[AB = \sqrt{(x_1-x_2)^2 + (y_1-y_2)^2}\] Any old memories resurfacing? :P

OpenStudy (anonymous):

ohhhh yeah. Damn. Do I have to do that 6 times D:

OpenStudy (anonymous):

gr I have to get up and get a pen and paper and turn my light on xD

OpenStudy (anonymous):

i think they are neither congruent nor similar. the ratios AB/PQ and BC/QR are not equal.

OpenStudy (anonymous):

Alright here goes nothing (everything) xD 1. (1,4)(-3,0) = (-2, -4) 2. (-2,-1)(-2,-4) = (-4,-5) 3. (3,-2)(2,3) = (1,-5) (-2,-4)(-4,-5)(1,-5) Did I do that right?

OpenStudy (anonymous):

I don't think I did. Damn it. Dx

OpenStudy (anonymous):

yea, u didn't. use the distance formula apoorvk showed above. and find out the lengths of any 2 corresponding sides. and then take their ratio. like, i found out AB and BC in triangle ABC and PQ and QR in triangle PQR. and then i took the ratio of the corr sides: AB/PQ and BC/QR.

OpenStudy (anonymous):

I"m screwing something up...Idk what though like the first set for example (1,4) (-3,0) -2^2 + -4^2 = 4 + 16 = 20 or something. idk D: Can one of you help me with just one of them and I can do the rest I swear.

OpenStudy (anonymous):

okay. lets find out AB first, then. A = (1,4) B = (-2,-1) so AB = \[\sqrt{(1 + 2)^{2} + (4+1)^{2}}\] \[= \sqrt{3^{2} + 5^{2}}\] \[= \sqrt{9+25}\] \[= \sqrt{34}\] \[= 5.83\]

OpenStudy (anonymous):

Okay! Thank you so much! Give me a few minutes for this

OpenStudy (anonymous):

sure, take ur time!

OpenStudy (anonymous):

Okay and I have: (1,4)(-2,-1) = √34 = 5.83 (-2,-1)(3,-2) = √34 = 5.83 (1,4)(3,-2) = √8 = 2.82 (-3,0)(-2,-4) = √51 = 7.14 (-2,-4)(2,3) = √65 = 8.06 (-3,0)(2,3) = √34 = 5.83 Is this good?

OpenStudy (anonymous):

Oh wait, (-3,0)(-2,-4) = √41 = 6.4

OpenStudy (anonymous):

(-2,-1)(3,-2) = √34 = 5.83 <=..........this should be sqrt(26) not of 34. i didn't confirm the others yet. but, u'd better re-check ur calculations.

OpenStudy (anonymous):

Is (1,4)(3,-2) = √40 = 6.32?

OpenStudy (anonymous):

yup. that's right.

OpenStudy (anonymous):

(1,4)(-2,-1) = √34 = 5.83 (-2,-1)(3,-2) = √26 = 5.1 (1,4)(3,-2) = √40 = 6.32 (-3,0)(-2,-4) = √41 = 6.4 (-2,-4)(2,3) = √65 = 8.06 (-3,0)(2,3) = √34 = 5.83 Okay. I think I got it. :D

OpenStudy (anonymous):

that's great! now just take any 2 ratios and see if their values match.

OpenStudy (anonymous):

ratios of corresponding sides, though!

OpenStudy (anonymous):

Well the very first and last ones do, I think? Or are those just angles?

OpenStudy (anonymous):

i didn't get u. u mean 5.83/6.4 and 6.32/5.83?

OpenStudy (anonymous):

I thought the 5.83 stand by themselves? I didn't know they become fractions :/

OpenStudy (anonymous):

I think AC and PR are similar?

OpenStudy (anonymous):

see, 2 triangles are similar if the ratios of their corresponding sides are equal. and they are congruent if their 'corresponding' sides are equal. here, we can clearly see that the latter is not the case. so we're seeing if they are similar by taking the ratios. got it?

OpenStudy (anonymous):

Okay, I think so. I'm pretty sure I do. The only similar sides are AC and PR, I think

OpenStudy (anonymous):

when u're testing the similarity between 2 triangles, u always take sides in pairs. so u can't have AC = PR and hence the 2 triangles are similar. as i said abv, take the ratio of corr sides. ie, AB-PQ, BC-QR, AC-PR. and check if these 3 ratios are equal or not.

OpenStudy (anonymous):

no.no.yes. ? so it is neither similar nor congruent?

OpenStudy (anonymous):

the answer should be just no. where did that yes come from?

OpenStudy (anonymous):

and yes, its neither similar nor congruent.

OpenStudy (anonymous):

oh .-. I'm really sorry if I'm frustrating! If I could give you like 100 medals I would. Thank you so much for helping me!

OpenStudy (anonymous):

lol! don't worry! i'm happy i could help u out! i hope u understood this and will be able to solve similar kinds in future. :)

OpenStudy (anonymous):

Thank you, thank you, thank you! (I can finally go to bed now :D) lol

OpenStudy (anonymous):

lol! kk...g'night!

OpenStudy (anonymous):

Night! :D